find the equation of each line in teh form y=mx+b if possible:
the line through (3,2) with undefined slope

Found 4 solutions by stanbon, Mathtut, jim_thompson5910, Alan3354:
Answer by stanbon(75887)

Equation: x = 3
=======================
Cheers,
Stan H.

Answer by Mathtut(3670)

point slope form takes on y-k=m(x-h) where (h,k) is a point on the line and m is the slope
:
so y-2=m(x-3)
:
y-2=mx-3m
:
y=mx-(3m-2) where m is the slope(undefined) and the y intercept is 3m-2

Answer by jim_thompson5910(35256)

the line through (3,2) with undefined slope is simply the equation x=3


Note: undefined slope equations are of the form x=a in which are vertical lines through points with the x coordinate of "a"

Answer by Alan3354(69443)

find the equation of each line in teh form y=mx+b if possible:
the line through (3,2) with undefined slope
---------------
y-y1 = m*(x-x1) (x1,y1) is (3,2)
y-2 = m*(x-3)
y = m*(x-3) + 2
Or, y = mx -3m +2